The function y(z, t)-(0.150 m) cos (та-15nt) where æ is in meters, and t is in seconds, describes a transverse wave on a taught string. a) What is the transverse velocity of a point on the string at an instant in time when y- +0.120 m? b) What is the maximum transverse velocity a point on the string can have? c) What is the average transverse velocity of a point on the string averaged over an integer number of wave periods?
d) What is the wave speed? e) How does the wave speed compare with the value you calculated in b)? Is there any meaningful comparison to be made between your answer to b) and e)?
Answer #1

y = 0.15 cos(pi x - 15 pi t)

at y = 0.12 ;

0.12 = 0.15 cos@

@ = 36.87 deg

Vy = 0.15 x 15 pi sin (pi x - 15 pi t)

= 0.15 x 15 pi x sin@

= 4.24 m/s

Vmax = 0.15 x 15 pi = 7.065 m/s

Vavg = 0 ... because wave

V = wavelength / T

but wavelength = 2 and T = 2/15 ... comparing with standard eqn

V = 2/(2/15) = 15 m/s

No comparison. wave speed is constant. Transverse speed changes with time and position